Samuel purchases a new car and finances $29,000 with a 5.9% loan over 6 years. Assuming each payment is due on the first of each month, what is the amount of Adam’s monthly car payment?
Answer #1

PV of ordinary annuity is mathematically represented as:

1- (1 + i)- PV of an Annuity Due =R*- i *(1 + i)

R = 5.9%/12 = 0.4917%

n = 6 * 12 = 72 months

90000D 1 -(1 + 0.004917)-72 . *(1 +0.004917) 0.004917

29000 = R * 60.51172 * 1.004917

R = $476.90